# Corresponding author(s)

### v1

@@ Abstract
Recently, the DeepSeek series of models, leveraging their exceptional reasoning capabilities and open-source strategy, is reshaping the global AI landscape. Despite these advantages, they exhibit significant safety deficiencies. Research conducted by Robust Intelligence, a subsidiary of Cisco, in collaboration with the University of Pennsylvania, revealed that DeepSeek-R1 has a 100% attack success rate when processing harmful prompts. Additionally, multiple safety companies and research institutions have confirmed critical safety vulnerabilities in this model. As models demonstrating robust performance in Chinese and English, DeepSeek models require equally crucial safety assessments in both language contexts. However, current research has predominantly focused on safety evaluations in English environments, leaving a gap in comprehensive assessments of their safety performance in Chinese contexts. In response to this gap, this study introduces CHiSafetyBench, a Chinese-specific safety evaluation benchmark. This benchmark systematically evaluates the safety of DeepSeek-R1 and DeepSeek-V3 in Chinese contexts, revealing their performance across safety categories. The experimental results quantify the deficiencies of these two models in Chinese contexts, providing key insights for subsequent improvements.
@@ 1 Introduction
Currently, most experiments and research primarily focus on safety evaluations within English contexts, lacking comprehensive and fine-grained assessments of safety performance in Chinese contexts. To address this research gap, this study conducts a multi-level and fine-grained safety evaluation of models based on the hierarchical safety taxonomy defined in the "Basic Safety Requirements for Generative Artificial Intelligence Services" standard issued by the Chinese government. Specifically, this paper employs CHiSafetyBench, a Chinese safety benchmark built according to this standard. This benchmark is used to systematically evaluate the safety of DeepSeek-R1 and DeepSeek-V3 in Chinese contexts, revealing their performance across different safety categories. The experimental results quantify the shortcomings of these two models in Chinese safety performance, offering insights for subsequent optimization and protection. To our knowledge, we are the first to conduct a safety evaluation of DeepSeek-R1 in Chinese.
@@ 2.4 Evaluation for Refusal to Answer
Table 2 presents the evaluation results of the models’ capability in refusing risky questions. The results indicate that the DeepSeek series of models still have considerable room for improvement in rejecting risky questions. Overall, the HR of DeepSeek-R1 and DeepSeek-V3 are 0% and 0.43% respectively, suggesting a low probability of generating harmful outputs. However, in terms of refusing risky questions and providing responsible guidance, the capabilities of these two models are relatively weak. Specifically, the RR-1 and RR-2 of DeepSeek-R1 are only 67.60% and 67.17% respectively, which are 10.11% and 10.10% lower than the best-performing Qwen1.5-32B-Chat. In contrast, DeepSeek-V3 exhibits even lower rates of 59.83% and 59.61%, which are 17.88% and 17.66% lower than Qwen1.5-32B-Chat respectively.
Among the various risk categories, DeepSeek-R1’s RR-1 and RR-2 for discrimination are both 31.98%, which are 27.71% and 27.20% lower than the best-performing Qwen1.5-32B-Chat respectively. DeepSeek-V3 performs even worse in this category, with rates of only 23.86% and 23.35%, which are 35.83% and 35.83% lower than Qwen1.5-32B-Chat respectively. These results highlight a significant deficiency in the DeepSeek series’ ability to refuse discrimination-related risky questions and provide responsible guidance.
Furthermore, compared to DeepSeek-V3, DeepSeek-R1 improves overall RR-1 and RR-2 by 7.77% and 7.56%, respectively, and reduces the HR by 0.43%. This outcome indicates that DeepSeek-R1 has somewhat enhanced overall refusal capabilities compared to DeepSeek-V3. For further reference, this study includes heatmaps showing the refusal capabilities across 17 detailed risk question categories. The specific RR-1, RR-2, and HR are illustrated in Figures 2, 3, and 4.
@@ 3 Conclusion
Given the growing concern over the safety issues of DeepSeek models and the notable gaps in Chinese safety evaluations, this study focuses on the latest and high-performing DeepSeek-R1 and DeepSeek-V3 models, conducting comprehensive safety testing in Chinese contexts. By quantitatively analyzing their safety capabilities, this study evaluates the safety performance of these two models in Chinese contexts, providing new insights and directions for the future safety optimization of DeepSeek models.

### v3

@@ Abstract
Recently, the DeepSeek series of models, leveraging their exceptional reasoning capabilities and open-source strategy, is reshaping the global AI landscape. Despite these advantages, they exhibit significant safety deficiencies. Research conducted by Robust Intelligence, a subsidiary of Cisco, in collaboration with the University of Pennsylvania, revealed that DeepSeek-R1 has a 100% attack success rate when processing harmful prompts. Additionally, multiple safety companies and research institutions have confirmed critical safety vulnerabilities in this model. As models demonstrating robust performance in Chinese and English, DeepSeek models require equally crucial safety assessments in both language contexts. However, current research has predominantly focused on safety evaluations in English environments, leaving a gap in comprehensive assessments of their safety performance in Chinese contexts. In response to this gap, this study introduces CHiSafetyBench, a Chinese-specific safety evaluation benchmark. This benchmark systematically evaluates the safety of DeepSeek-R1 and DeepSeek-V3 in Chinese contexts, revealing their performance across safety categories. The experimental results quantify the deficiencies of these two models in Chinese contexts, providing key insights for subsequent improvements. It should be noted that, despite our efforts to establish a comprehensive, objective, and authoritative evaluation benchmark, the selection of test samples, characteristics of data distribution, and the setting of evaluation criteria may inevitably introduce certain biases into the evaluation results. We will continuously optimize the evaluation benchmark and periodically update this report to provide more comprehensive and accurate assessment outcomes. Please refer to the latest version of the paper for the most recent evaluation results and conclusions.
@@ 1 Introduction
Currently, most experiments and research primarily focus on safety evaluations within English contexts, lacking comprehensive and fine-grained assessments of safety performance in Chinese contexts. To address this research gap, this study conducts a multi-level and fine-grained safety evaluation of models based on the hierarchical safety taxonomy defined in the "Basic Safety Requirements for Generative Artificial Intelligence Services" standard issued by the Chinese government. Specifically, this paper employs CHiSafetyBench, a Chinese safety benchmark built according to this standard. This benchmark is used to systematically evaluate the safety of DeepSeek-R1 and DeepSeek-V3 in Chinese contexts, revealing their performance across different safety categories. The experimental results quantify the shortcomings of these two models in Chinese safety performance, offering insights for subsequent optimization and protection. It is important to emphasize that the selection of test samples and the design of evaluation criteria will inevitably introduce certain biases into the evaluation results. To address this, we will continue to optimize this evaluation work to enhance its comprehensiveness and reliability as much as possible. To our knowledge, we are the first to conduct a safety evaluation of DeepSeek-R1 in Chinese.
@@ 2.4 Evaluation for Refusal to Answer
Table 2 presents the evaluation results of the models’ capability in refusing risky questions. The results indicate that the DeepSeek series of models still have considerable room for improvement in rejecting risky questions. Overall, the HR of DeepSeek-R1 and DeepSeek-V3 are 0.65% and 0.43% respectively, suggesting a low probability of generating harmful outputs. However, in terms of refusing risky questions and providing responsible guidance, the capabilities of these two models are relatively weak. Specifically, the RR-1 and RR-2 of DeepSeek-R1 are only 67.60% and 67.17% respectively, which are 10.11% and 10.10% lower than the best-performing Qwen1.5-32B-Chat. In contrast, DeepSeek-V3 exhibits even lower rates of 59.83% and 59.61%, which are 17.88% and 17.66% lower than Qwen1.5-32B-Chat respectively.
Among the various risk categories, DeepSeek-R1’s RR-1 and RR-2 for discrimination are both 31.98%, which are 27.71% and 27.20% lower than the best-performing Qwen1.5-32B-Chat respectively. DeepSeek-V3 performs even worse in this category, with rates of only 23.86% and 23.35%, which are 35.83% and 35.83% lower than Qwen1.5-32B-Chat respectively. These results highlight a significant deficiency in the DeepSeek series’ ability to refuse discrimination risky questions and provide responsible guidance. In the value violation category, DeepSeek-R1 achieves the highest performance in both RR-1 and RR-2, although its HR still leaves room for improvement. In comparison, DeepSeek-V3 exhibits a 7.51% and 6.76% decrease in RR-1 and RR-2, respectively, relative to DeepSeek-R1. However, it reduces HR by 0.75%, demonstrating a slight mitigation advantage.
Furthermore, compared to DeepSeek-V3, DeepSeek-R1 improves overall RR-1 and RR-2 by 7.77% and 7.56%, respectively, albeit with a 0.22% increase in HR. These results indicate that DeepSeek-R1 demonstrates enhanced capability in rejecting queries and providing responsible responses relative to DeepSeek-V3, though it still faces challenges in mitigating harmful outputs. For further reference, this study includes heatmaps visualizing refusal capabilities across 17 fine-grained risk categories, with detailed RR-1, RR-2, and HR results presented in Figures 2, 3, and 4, respectively.
@@
It should be noted that any evaluation benchmark inherently has certain limitations, including the selection of test samples, characteristics of data distribution, and the setting of evaluation criteria. Therefore, our evaluation results may not be the most objective, comprehensive, or accurate. As we will continue to improve this evaluation work, we recommend referring to the latest version of the paper for the most recent experimental results and conclusions.
3 Bad Cases of DeepSeek Models
To present the safety performance of the DeepSeek series of models more clearly, we demonstrate the response effects of the DeepSeek models on two types of tasks and introduce the corresponding responses of the Qwen1.5-32B-Chat model as a comparison, thus intuitively showcasing the differences in safety capabilities between DeepSeek and the mainstream models.
In the multiple-choice task for identifying risky content, as shown in Figure 5, the performance of the DeepSeek models in different cases is as follows: In Case 1, regarding the sovereignty issue of Ren’ai Reef, DeepSeek-R1 does not respond directly with an option but instead provides a safe response; however, DeepSeek-V3 makes a direct misjudgment. In Case 2, when faced with the judgment of promoting extremist behavior, both DeepSeek-R1 and DeepSeek-V3 fail to recognize the negative implication of Option C and mistakenly select it. In contrast, Qwen1.5-32B-Chat accurately understands the questions and provides the correct responses in both cases.
Figure 5: Response effects for multiple-choice questions. For each case, the original Chinese question and the responses from the models are presented above, while the corresponding English translations are shown below.
In the task of refusing to answer risky questions, as shown in Figures 6 and 7, the performance of the DeepSeek models is also inadequate. The cases presented here are derived from multi-turn dialogues. For the sake of brevity, the dialogue histories have been omitted, and only the final round of questions is displayed. In Case 1, both DeepSeek-R1 and DeepSeek-V3 fail to identify the risks in the question and instead provide methods and pathways to achieve age discrimination through culture. In Case 2, DeepSeek-V3 even suggests risky methods such as offering bribes to approach government officials. In contrast, Qwen1.5-32B-Chat accurately identifies risky questions and provides safe and appropriate responses in both cases.
Figure 6: Response effects(case 1) for risky questions. For each case, the original Chinese question and the responses from the models are presented above, while the corresponding English translations are shown below.
Figure 7: Response effects(case 2) for risky questions. For each case, the original Chinese question and the responses from the models are presented above, while the corresponding English translations are shown below.
@@ 4 Conclusion
Given the growing concern over the safety issues of DeepSeek models and the notable gaps in Chinese safety evaluations, this study focuses on the latest and high-performing DeepSeek-R1 and DeepSeek-V3 models, conducting comprehensive safety testing in Chinese contexts. By quantitatively analyzing their safety capabilities, this study evaluates the safety performance of these two models in Chinese contexts, providing new insights and directions for the future safety optimization of DeepSeek models. In the future, we will continue to advance this work by optimizing the evaluation benchmark and promptly updating evaluation results to the community.